# Evaluating the Elementary Multilingual Capabilities of Large Language Models with MultiQ

### v1

@@ 4 Drivers of Multilinguality
Both Llama2 and Mistral employ a tokenizer that uses byte pair encoding (BPE) (Sennrich et al., 2016), streamlining sequence encoding by minimizing token count through identifying common subwords. This allows for frequent terms to be represented through subwords, whereas rare or unseen terms prompt the model to default to individual characters or ultimately ASCII codes for representation.
@@
First, we evaluate how different tokenization strategies impact prompt representation in MultiQ. We observe a significant deviation in the number of unique tokens used to represent all prompts in MultiQ: Mistral uses 9,933 unique tokens, contrasting with Llama2’s 10,676. This suggests that Llama2 may be less efficient at segmenting the typologically diverse MultiQ dataset into a smaller number of subwords.
@@
For each model, we group languages into three categories depending on the model’s tokenization strategy fo the respective language. We develop a heuristic that allows us to classify languages into three tokenization categories: “ASCII”, “character” or “subword”. We do so by compiling the model’s 20 most commonly used tokens for representing each language in MultiQ, removing noise tokens (e.g., sentence start) from the list, and then quantifying the prevalence of ASCII tokens and characters based on the token-id ranges they usually occupy. Languages with over 70% of tokens in the ASCII or character categories were classified accordingly. For Llama2, out of 137 languages, we find 89 subword, 36 character and 12 ASCII languages. Mistral has 37 in character and 9 in ASCII. The models differ in the tokenization of some symbolic languages such as Chinese (zh) (see Figure 6), compared to Mistral mostly using individual characters, Llama2 more frequently has to use ASCII tokens due to its limited recognition of Chinese tokens, hindering its ability to effectively map Chinese vocabulary.
@@
We evaluate the models’ average QA accuracy across tokenization categories, finding a clear hierarchy (see Table 5): Subword encoding outperforms character and ASCII encodings, with Mistral achieving 20.4% accuracy on subword encoded languages and Llama2 24%. Acknowledging the limitations of our heuristic in precisely classifying tokenization strategies, we believe that exploring tokenization optimization is a promising direction for multilingual research, which also aligns with the broader discourse on the impact of tokenization on model capabilities. We hope these insights highlight the importance of alternative language representation strategies (e.g., pixel-based models (Salesky et al., 2023)).
Figure 6: Tokenization analysis. Distribution of the most used unique tokens to represent all prompts in Chinese (zh), given the respective tokenizers of the Llama2 (7B) and Mistral (7B).
@@
Table 5: QA accuracy on MultiQ split by tokenization strategy. Highest accuracy per model is bold.
@@ 5 Related Work
Existing multilingual benchmarks primarily target the performance of fully or partially supervised models on collections of standard NLP tasks, like XNLI (Conneau et al., 2018). Popular examples are the XTREME benchmarks (e.g., Ruder et al., 2021, 2023), and XGLUE (Liang et al., 2020). Similarly, researchers also presented benchmarks designed for languages spoken in particular regions, e.g., IndicXTREME (Doddapaneni et al., 2023) for Indic languages, Masakhan-NER (Adelani et al., 2022), and TaTA (Gehrmann et al., 2023) covering African languages, as well as NusaX (Winata et al., 2023) for Indonesian languages. Recently, Ahuja et al. (2023) proposed MEGA for evaluating multilingual generative models, which they use to evaluate LLMs like GPT-4 on a set of standard tasks. In a similar vein, Asai et al. (2023) presented BUFFET and benchmark LLMs for few-shot transfer. Artetxe et al. (2020) present XQuaD, a multilingual QA dataset focusing on reading comprehension. All these benchmarks typically cover roughly 20 to 90 languages. In contrast, MultiQ contains parallel questions in 137 languages covering 95.38% of Grambank features, which allows us to study the multilingual behavior of LLMs at the margins.
Multilingual LLMs.
In the past, researchers have presented a range of multilingual LMs like mBERT (Devlin et al., 2019), XLM-R (Conneau et al., 2020), and Glot-500 ImaniGooghari et al. (2023). These models are most effective in a supervised fine-tuning setting. Different from that, models like mT5 (Xue et al., 2021), BLOOMZ (Workshop et al., 2022), Okapi (Lai et al., 2023b) and the very recently released Aya (Üstün et al., 2024) better allow for prompting. For all those models, many studies also explored how to further adapt them to more languages (e.g., Yong et al., 2023a), assessed their detailed capabilities (e.g., Zhang et al., 2023a; Yong et al., 2023b), and aspired to understand the cross-lingual mechanisms underpinning their capabilities (e.g., Lauscher et al., 2020; Choenni et al., 2023, inter alia). Given the scale of large chat models, the landscape of open models that cover multiple languages is, however, still scarce.
Multilinguality in Monolingual LLMs.
@@
Given that multilingual chat models are barely openly available, we are especially interested in assessing the multilinguality of supposedly English-only models. Blevins and Zettlemoyer (2022) explained this behavior through data contamination: while the vast majority of the pre-training data of those models is English (e.g., ~93% for GPT-3 (Brown et al., 2020), and ~90% for Llama-2 (Touvron et al., 2023a)) there are also small portions of non-English training data present. In such cases, the dominant language can help unlock the models’ capabilities for the underrepresented languages (Gogoulou et al., 2022). As a result, recent research assesses the multilinguality of several English-centric models, like GPT-3 and ChatGPT (e.g., Zhang et al., 2023b; Lai et al., 2023a; Armengol-Estapé et al., 2022; Winata et al., 2021). However, all these works focus on standard NLP tasks, one or few models, and on a few major languages only, while we analyze the multilingual behavior of six LLMs at scale.
@@ 6 Conclusion
We introduced MultiQ, a new silver standard benchmark for open-ended question answering that covers 137 typologically diverse languages. With MultiQ, we evaluated the basic multilingual capabilities of six current, chat-optimized open LLMs, which are not all intended for multilingual use. Our analysis focused on two key dimensions of multilingual capability – language fidelity and QA accuracy – and how they relate to each other. We found that some models designed primarily for English reveal significant capabilities in one dimension or the other. We noticed that although language fidelity is not decisive, it positively impacts QA accuracy, prompting us to explore other underlying factors. We hope that our work will spark further research in the direction of the multilingual capability of open chat-optimized models, with a particular emphasis on diverse multilingual dimensions.
@@
Our work comes with several limitations that have already been partially discussed throughout the paper. First, we automatically translate our dataset, which introduces noise in the dataset, however, given our detailed validation, we strongly believe that the implications of our findings still hold. Second, the automated evaluation of the models’ answering accuracy using GPT-4 and their response language using GlotLID are additional factors that might bring noise into the results. However, further validation assures us that these factors minimally impact the results, potentially leading to an underestimation of answer accuracy. Third, mapping Google Translate IDs to the ISO language codes of GlotLID wasn’t always directly possible, but we excluded possible inaccuracies from our calculations. Lastly, our analysis intentionally concentrated on basic multilingual capabilities, excluding the assessment of advanced reasoning or formulation skills.
@@ Appendix A MultiQ
Code
Language
Family Code
Language
Family Code
Language
Family
af
Afrikaans
IE hmn
Hmong
HM ny
Nyanja (Chichewa)
NC
ak
Twi (Akan)
NC hr
Croatian
IE om
Oromo
AA
am
Amharic
AA ht
Haitian Creole
or
Odia (Oriya)
IE
ar
Arabic
AA hu
Hungarian
U pa
Panjabi
IE
as
Assamese
IE hy
Armenian
IE pl
Polish
IE
ay
Aymara
Ay id
Indonesian
Au ps
Pashto
IE
az
Azerbaijani
Al ig
Igbo
NC pt
Portuguese
IE
be
Belarusian
IE ilo
Ilocano
Au qu
Quechua
Qu
bg
Bulgarian
IE is
Icelandic
IE ro
Romanian
IE
bho
Bhojpuri
IE it
Italian
IE ru
Russian
IE
bm
Bambara
M iw
Hebrew alternativ
AA rw
Kinyarwanda
NC
bn
Bengali
IE ja
Japanese
J sa
Sanskrit
bs
Bosnian
IE jv
Javanese alternativ
Au sd
Sindhi
IE
ca
Catalan
IE jw
Javanese
Au si
Sinhala (Sinhalese)
IE
ceb
Cebuano
Au ka
Georgian
K sk
Slovak
IE
ckb
Kurdish (Sorani)
IE $k$
Kazakh
Al sl
Slovenian
IE
co
Corsican
IE km
Khmer
AuA sm
Samoan
Au
cs
Czech
IE kn
Kannada
D sn
Shona
NC
cy
Welsh
IE ko
Korean
K so
Somali
AA
da
Danish
IE kri
Krio
sq
Albanian
IE
de
German
IE ku
Kurdish
IE sr
Serbian
IE
doi
Dogri
IE ky
Kyrgyz
Al st
Sesotho
NC
dv
Dhivehi
IE la
Latin
su
Sundanese
Au
$e$
Ewe
NC lb
Luxembourgish
IE sv
Swedish
IE
el
Greek
IE lg
Luganda
NC sw
Swahili
NC
en
English
IE ln
Lingala
NC ta
Tamil
D
eo
Esperanto
IE lo
Lao
TK te
Telugu
D
es
Spanish
IE lt
Lithuanian
IE tg
Tajik
IE
et
Estonian
U lus
Mizo
ST th
Thai
TK
eu
Basque
B lv
Latvian
IE ti
Tigrinya
AA
fa
Persian
IE mai
Maithili
IE tk
Turkmen
Al
fi
Finnish
U mg
Malagasy
Au tl
Tagalog (Filipino)
Au
fil
Filipino (Tagalog)
Au mi
Maori
Au tr
Turkish
Al
fr
French
IE mk
Macedonian
IE ts
Tsonga
NC
fy
Frisian
IE ml
Malayalam
D $t$
Tatar
Al
ga
Irish
IE mn
Mongolian
ug
Uyghur
Al
gd
Scots Gaelic
IE mni-Mtei
Meiteilon (Manipuri)
ST uk
Ukrainian
IE
gl
Galician
IE mr
Marathi
IE ur
Urdu
IE
gn
Guarani
T ms
Malay
Au uz
Uzbek
Al
gom
Konkani
IE mt
Maltese
AA vi
Vietnamese
AuA
gu
Gujarati
IE my
Myanmar (Burmese)
ST xh
Xhosa
NC
ha
Hausa
AA ne
Nepali
IE yi
Yiddish
IE
haw
Hawaiian
Au nl
Dutch
IE yo
Yoruba
NC
he
Hebrew
AA no
Norwegian
IE zh
Chinese (Trad.)
hi
Hindi
IE nso
Sepedi
NC zh-CN
Chinese (Simpl.)
zh-TW
Chinese (Simpl.)
zu
Zulu
NC
@@ Appendix B Experimental Setup
Task
Prompt
Question Creation for MultiQ
Generate a list of 10 simple and clear {domain} questions with simple and clear answers.
Validate Answer Correctness
For the following question, is this a correct answer?
@@
Answer: {Model Answer}
The question is in English. The answer is either in English or in another language. If the answer is correct, type ’yes’. If the answer is incorrect, type ’no’."

### v2

@@ 4 Tokenization and Multilinguality
Both Llama2 and Mistral employ a tokenizer that uses byte pair encoding (BPE, Sennrich et al., 2016), streamlining sequence encoding by minimizing token count through identifying common subwords. This allows for frequent sequences to be represented through subword tokens, whereas rare or unseen sequences make the model default to individual characters or, failing that, ASCII code tokens.
@@
First, we evaluate how different tokenization strategies impact prompt tokenization in MultiQ. We observe a significant deviation in the number of unique tokens used to represent all prompts in MultiQ: Mistral uses 9,933 unique tokens, contrasting with Llama2’s 10,676. This suggests that Llama2 may be less efficient at segmenting the typologically diverse MultiQ dataset into a smaller number of subwords.
@@
For each model, we group languages into three categories depending on the model’s tokenization strategy for the respective language. We develop a heuristic that allows us to classify languages into three tokenization categories: “ASCII”, “character” or “subword”. We do so by compiling the model’s 20 most commonly used tokens for representing each language in MultiQ, removing noise tokens (e.g. sentence start) from the list, and then quantifying the prevalence of ASCII tokens and characters based on the token-id ranges they usually occupy. Languages with over 70% of tokens in the ASCII or character categories were classified accordingly. For Llama2, out of 137 languages, we find 89 subword, 36 character and 12 ASCII languages. Mistral has 37 character and 9 ASCII languages. The models differ in the tokenization of some symbolic languages such as Chinese (zh, see Figure 6): while Mistral mostly uses individual character tokens, Llama2 more frequently resorts to ASCII tokens due to its limited Chinese token vocabulary, which limits its ability to effectively tokenize Chinese language prompts.
@@
We evaluate the models’ average QA accuracy across tokenization categories, finding a clear hierarchy (see Table 5): Subword encoding outperforms character and ASCII encodings, with Mistral and Llama2 respectively achieving 20.4% and 24.0% accuracy on subword-encoded languages compared to just 6.4% and 11.4% on character-encoded languages. Even though our heuristic may introduce some noise in classifying tokenization strategies, we believe that exploring tokenization optimization is a promising direction for multilingual research, which also aligns with the broader discourse on the impact of tokenization on model capabilities. We hope our insights can help motivate further research into alternative language representation strategies (e.g. pixel-based models, Salesky et al., 2023).
Figure 6: Tokenization analysis. Distribution of the most used unique tokens across all n=200 MultiQ prompts in Chinese (zh), when using the tokenizers of Llama2 (7B) and Mistral (7B).
@@
Table 5: QA accuracy (%) on MultiQ split by tokenization strategy. Highest accuracy per model is bold.
@@ 5 Related Work
Existing multilingual benchmarks primarily target the performance of fully or partially supervised models on collections of standard NLP tasks, like XNLI (Conneau et al., 2018). Popular examples are the XTREME benchmarks (e.g. Ruder et al., 2021, 2023), and XGLUE (Liang et al., 2020). Similarly, researchers also presented benchmarks designed for languages spoken in particular regions, e.g. IndicXTREME (Doddapaneni et al., 2023) for Indic languages, Masakhan-NER (Adelani et al., 2022), and TaTA (Gehrmann et al., 2023) covering African languages, as well as NusaX (Winata et al., 2023) for Indonesian languages. Recently, Ahuja et al. (2023) proposed MEGA for evaluating multilingual generative models, which they use to evaluate LLMs like GPT-4 on a set of standard tasks. In a similar vein, Asai et al. (2023) presented BUFFET and benchmark LLMs for few-shot transfer. Another line of work focuses on multilingual QA datasets for reading comprehension, such XQuaD Artetxe et al. (2020), TyDiQA (Clark et al., 2020) and the Belebele benchmark (Bandarkar et al., 2023) with up to 122 diverse languages. Concurrent work proposes a new multilingual instruction tuning dataset called Aya (Singh et al., 2024), which also covers a range of open-ended questions in 114 languages. By comparison, MultiQ contains parallel questions in a larger set of 137 languages covering 95.4% of Grambank features, which demonstrates its typological diversity and allows the analysis of multilingual LLM behavior at the margins of language coverage. Furthermore, MultiQ’s carefully selected short and simple questions target basic LLM knowledge to test only their multilingual capabilities and not ancillary factors such as complex reasoning.
Multilinguality in Monolingual LLMs
@@
Given the limited availability of open multilingual chat models, we are especially interested in assessing the multilinguality of models intended for English use only. Blevins and Zettlemoyer (2022) explained this behavior through data contamination: while the vast majority of the pre-training data of those models is English (e.g. ~93% for GPT-3 Brown et al., according to 2020, and ~90% for Llama-2 according to Touvron et al., 2023) there are also small portions of non-English content in the pre-training data. In such cases, it appears that the dominant language can help “unlock” the models’ capabilities for the underrepresented languages (Gogoulou et al., 2022). Consequently, recent research assesses the multilinguality of several English-centric models, like GPT-3 and ChatGPT (e.g. Zhang et al., 2023; Lai et al., 2023; Armengol-Estapé et al., 2022; Winata et al., 2021). However, these prior works focus on standard NLP tasks such as text classification, just one or few models, and on just a few major languages. By contrast, we test the multilingual behavior of six LLMs in a more natural open-ended QA setting.
@@ 6 Conclusion
We introduced MultiQ, a new silver standard benchmark for open-ended question answering that covers 137 typologically diverse languages. With MultiQ, we evaluated the basic multilingual capabilities of six current, chat-optimized open LLMs, which are restricted in their intended use to just one or a small handful of languages. Our analysis focused on two key dimensions of multilingual capability – language fidelity and QA accuracy – and how they relate to each other. We found that all LLMs we test respond faithfully and/or accurately for at least some languages beyond their intended use. Most models are more accurate when they respond faithfully. However, we found that differences across models are large, and that there is a long tail of languages where models are neither accurate nor faithful. Finally, we indentified differences in tokenization as a potential explanation for our results. Overall, we hope that our findings can motivate further research into improving the multilingual capabilities of open LLMs, especially for diverse and under-represented languages, so that language technologies can benefit everyone, regardless of which language they speak.
@@
Our work comes with several limitations, which have already been partially discussed throughout the paper. First, we automatically translate our dataset, which introduces noise in the dataset. However, based on manual validation by native speakers, translation quality is high. Second, the automated evaluation of the models’ answering accuracy using GPT-4, as well as response language classification using GlotLID, introduce some noise into the results. Here, too, human validation assures us that these factors minimally impact the results, potentially leading to an underestimation of answer accuracy. Third, mapping Google Translate IDs to the ISO language codes of GlotLID was not always directly possible, but we excluded possible inaccuracies from our calculations. Lastly, our analysis intentionally concentrated on basic multilingual capabilities, excluding the assessment of advanced reasoning or formulation skills.
@@ Appendix A MultiQ
Code Language Family Code Language Family Code Language Family
af Afrikaans IE hmn Hmong HM ny Nyanja (Chichewa) NC
ak Twi (Akan) NC hr Croatian IE om Oromo AA
am Amharic AA ht Haitian Creole or Odia (Oriya) IE
ar Arabic AA hu Hungarian U pa Panjabi IE
as Assamese IE hy Armenian IE pl Polish IE
ay Aymara Ay id Indonesian Au ps Pashto IE
az Azerbaijani Al ig Igbo NC pt Portuguese IE
be Belarusian IE ilo Ilocano Au qu Quechua Qu
bg Bulgarian IE is Icelandic IE ro Romanian IE
bho Bhojpuri IE it Italian IE ru Russian IE
bm Bambara M iw Hebrew alternativ AA rw Kinyarwanda NC
bn Bengali IE ja Japanese J sa Sanskrit
bs Bosnian IE jv Javanese alternativ Au sd Sindhi IE
ca Catalan IE jw Javanese Au si Sinhala (Sinhalese) IE
ceb Cebuano Au ka Georgian K sk Slovak IE
ckb Kurdish (Sorani) IE $k$ Kazakh Al sl Slovenian IE
co Corsican IE km Khmer AuA sm Samoan Au
cs Czech IE kn Kannada D sn Shona NC
cy Welsh IE ko Korean K so Somali AA
da Danish IE kri Krio sq Albanian IE
de German IE ku Kurdish IE sr Serbian IE
doi Dogri IE ky Kyrgyz Al st Sesotho NC
dv Dhivehi IE la Latin su Sundanese Au
$e$ Ewe NC lb Luxembourgish IE sv Swedish IE
el Greek IE lg Luganda NC sw Swahili NC
en English IE ln Lingala NC ta Tamil D
eo Esperanto IE lo Lao TK te Telugu D
es Spanish IE lt Lithuanian IE tg Tajik IE
et Estonian U lus Mizo ST th Thai TK
eu Basque B lv Latvian IE ti Tigrinya AA
fa Persian IE mai Maithili IE tk Turkmen Al
fi Finnish U mg Malagasy Au tl Tagalog (Filipino) Au
fil Filipino (Tagalog) Au mi Maori Au tr Turkish Al
fr French IE mk Macedonian IE ts Tsonga NC
fy Frisian IE ml Malayalam D $t$ Tatar Al
ga Irish IE mn Mongolian ug Uyghur Al
gd Scots Gaelic IE mni-Mtei Meiteilon (Manipuri) ST uk Ukrainian IE
gl Galician IE mr Marathi IE ur Urdu IE
gn Guarani T ms Malay Au uz Uzbek Al
gom Konkani IE mt Maltese AA vi Vietnamese AuA
gu Gujarati IE my Myanmar (Burmese) ST xh Xhosa NC
ha Hausa AA ne Nepali IE yi Yiddish IE
haw Hawaiian Au nl Dutch IE yo Yoruba NC
he Hebrew AA no Norwegian IE zh Chinese (Trad.)
hi Hindi IE nso Sepedi NC zh-CN Chinese (Simpl.)
zh-TW Chinese (Simpl.) zu Zulu NC
@@ Appendix B Experimental Setup
Task Prompt
Question Creation for MultiQ Generate a list of 10 simple and clear {domain} questions with simple and clear answers.
Validate Answer Correctness For the following question, is this a correct answer?
@@
Answer: {Model Answer} The question is in English. The answer is either in English or in another language. If the answer is correct, type ’yes’. If the answer is incorrect, type ’no’."
@@ Appendix C Granular Language Fidelity Analysis
Prompt Language English Translation Model Completion
Xana Misava yi na tin’hweti tingani? ts How many moons does Earth have? Xana Misava yi na tin’hweti tingani? (What is Xana Misava’s net worth?)
X’inhu sena dawl? mt What is a light-year? X’inhu sena dawl? is a question in the Mayan language of Guatemala. It translates to "What is the meaning of life?" in English
Beth yw mynydd uchaf Japan? cy What is the highest mountain in Japan? Beth yw mynydd uchaf Japan? (Do you speak Japanese?)
Quanto fa tre volte sette? it What is three times seven? Quanto fa tre volte sette? Quanto fa tre volte sette? Quanto fa tre volte sette? Quanto fa tre volte sette? Quanto fa tre volte sette? Quanto fa tre volte sette? Quanto fa tre volte sette? Quanto fa tre volte sette? Quanto fa tre volte sette? Quanto fa tre volte sette? […]
Table 11: Examples of Fidelity Issues for Mistral 7B
@@ Appendix D Answer Accuracy
Appendix E Correlation between Answer Accuracy and Language Fidelity